\begin{thebibliography}{99}




%\bibitem{Behzad}
%M. Behzad, G. Chartrand and C.E. Wall, On minimal regular digraphs
%with given girth, Fund. Math. 69 (1970), 227-231.
\bibitem{Bondy}
J. A. Bondy, Counting subgraphs: a new approach to the
Caccetta-H\"aggkvist conjecture. Graphs and combinatorics
(Marseille, 1995). Discrete Math. 165/166 (1997), 71-80.

 \bibitem{Bondy2} J. A. Bondy,  U. S. R. Murty,
Graph theory.
Graduate Texts in Mathematics, 244. Springer, New York, 2008.
\bibitem{CH} L. Caccetta and R. H\"aggkvist, On minimal  digraphs with given girth,  Proceedings of the Ninth Southeastern Conference on Combinatorics, Graph Theory, and Computing (Florida Atlantic Univ., Boca Raton, Fla., 1978), (Winnipeg, Man.),  Congress. Numer., XXI, Utilitas Math.  (1978), 181-187.

\bibitem{el} S. Eliahou and C. Lecouvey, Matchings in arbitrary groups, {\it Advances in Applied Mathematics} 40 (2008) 219–224.

\bibitem{halll} P. Hall, On representatives of subsets, {\it J. London Math. Soc.} 10 (1935),
26-30.

\bibitem{EH} P. Erd\H{o}s and H. Heilbronn,
On the Addition of residue classes mod  $p$, {\it Acta Arith.} {\bf
9} (1964), 149-159.
\bibitem{gaotnd} W. D. Gao,
A combinatorial problem on finite abelian groups.
{\it J. Number Theory} 58 (1996), no. 1, 100--103.

\bibitem{gerlodinger} A. Geroldinger, F. Halter-Koch,  Non-unique factorizations.
Algebraic, combinatorial and analytic theory. Pure and Applied
Mathematics (Boca Raton), 278. Chapman \& Hall/CRC, Boca Raton, FL,
2006. xxii+700 pp.

\bibitem{HATOM} Y.O. Hamidoune, Sur les atomes d'un graphe orient\'e,
 C.R. Acad. Sc. Paris A 284 (1977),   1253-1256.


\bibitem{HJCT} Y.O. Hamidoune,   Quelques probl\`emes de connexit\'e dans les
graphes orient\'es, J. Comb. Theory B 30 (1981), 1-10.


\bibitem{HEJC} Y.O. Hamidoune,   An application  of  connectivity theory in graphs to
      factorizations of elements in groups, Europ. J of Combinatorics 2 (1981), 349-355.

\bibitem{Hdm} Y.O. Hamidoune,   Sur les atomes d'un graphe de Cayley infini,
 Discrete Math., 73 (1989), 297-300.

 \bibitem{Hast} Y.O. Hamidoune,  On small subset product in a group.
Structure Theory of set-addition, Ast\'erisque. no. 258(1999),
xiv-xv, 281-308.

\bibitem{Hspheres}Y. O. Hamidoune. On iterated image size for point-symmetric
relations. Combin. Probab. Comput. 17 (2008), no. 1, 61--66.

\bibitem{HweightD}
 Y. O. Hamidoune, A weighted generalization of Gao's $n+D-1$--Theorem,
Combin. Probab. Comput. 17 (2008), no. 17, 793--798.

\bibitem{kempcompl} J.H.B. Kempermann, On complexes in a semigroup,
{\it Nederl. Akad. Wetensch. Proc. Ser. A.} 59  {\it Indag. Math.}
18 (1956), 247--254.

\bibitem{los} J. Losonczy, On matchings in groups, Adv. in Appl. Math. 20 (1998) 385–391.

\bibitem{mader} W. Mader, Existence of openly disjoint circuits through a vertex. Prprint received by e-mail from the author on January, 30-2008.


    \bibitem{Nat}M. B. Nathanson,
{\it Additive Number Theory. Inverse problems and the geometry of
sumsets}, Grad. Texts in Math. 165, Springer, 1996.

%\bibitem{lovasz} L. Lov\'asz, M.D. Plummer, Matching theory, Ann. Discrete Math. 29 (1986) 137.

\bibitem{Nat0} M. B. Nathanson, The Caccetta-H\"aggkvist conjecture
and Additive Number Theory, arXiv:math. CO/0603469, 2006.



\bibitem{olsonaa} J.E. Olson, Sums of sets of group elements,
{\it Acta Arith.} 28 (1975/76), no. 2, 147--156.

\bibitem{sch}{ P. Scherk}, {\em Distinct elements in a set of sums}, Amer. Math. Monthly, 62 (1955), pp. 46--47.

\bibitem{sheph} J. C. Shepherdson, On the addition of elements of a sequence,  J. London Math Soc. 22(1947), 85-88.

    \bibitem{tv} T. Tao, V.H. Vu,  {\it Additive Combinatorics}, Cambridge Studies
in Advanced Mathematics 105 (2006), Cambridge University Press.


\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{amsmath}

\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{euscript}
\newenvironment{proof}{\noindent {\it Proof.~~}\ }{\  \rule{1mm}{2mm}\medskip}

\newenvironment{proof*}{\noindent {\it Proof.~~}\ }{}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{theirtheorem}{Theorem}

\def\G{\partial}

\newcommand{\oa}{\overline{A}}

\begin{document}
\title{Extensions of the  Moser-Scherck-Kemperman-Wehn Theorem }
\author{ {Y. O. Hamidoune}\thanks{
UPMC Univ Paris 06,
 E. Combinatoire, Case 189, 4 Place Jussieu,
75005 Paris, France.}
}

\maketitle

%11B60, 11B34, 20D60.

\begin{abstract} Let $\Gamma =(V,E)$  be a  reflexive
relation having a transitive group of automorphisms and  let $v\in V.$
Let $F$  be a  subset of $V$ with $F\cap \Gamma ^-(v)=\{v\}$.
\begin{itemize}
  \item[(i)] If $F$  is finite, then
$| \Gamma (F)\setminus F|\ge
|\Gamma (v)|-1.$
\item [(ii)]If $F$  is cofinite, then
$| \Gamma (F)\setminus F|\ge
|\Gamma ^- (v)|-1.$
\end{itemize}

In particular,
let $G$  be group, $B$ be a finite subset of $G$
and let $F$ be a finite or a cofinite subset of $G$ such that  $F\cap  B^{-1}=\{1\}$.
Then
$| (FB)\setminus F|\ge
|B|-1.$
The last result (for $F$ finite), is  famous  Moser-Scherck-Kemperman-Wehn Theorem. Its   extension
to cofinite subsets seems new. We give also few applications.

 \end{abstract}

\section{Introduction}

A problem of Moser solved also by Scherck in \cite{sch} states that in ablian group $G$ two finite subsets $A,B$ with $|A\cap B^{-1}|=\{1\}$ verifies the Cauchy-Davenport inequality: $|AB|\ge |A|+|B|-1.$ The validity of this result in the non-abelian case was proved by Kemperman \cite{kempcompl}, mentioning that the result  was independently proved
by When \cite{kempcompl}. In our work,   $A$ could be  infinite. So we shall use the formulation $|AB\setminus A|\ge |B|-1.$

All the known proofs of the Moser-Scherck-Kemperman-Wehn Theorem, used additive transforms. In this work, we obtain a completely different proof.
The Moser-Scherck-Kemperman-Wehn Theorem has important applications
in Number Theory and the reader may find several applications of this beautiful result to the Theory of Non-unique factorization in the text book of Geroldinger-Halter-Koch \cite{gerlodinger}. Recall that this result  is used among other tools by Olson
\cite{olsonaa},  to prove that a  subset $S$ of a finite group $G$ with $|S|\ge 3\sqrt{|G|}$ contains a non-empty subset with a product (under some ordering) $=1$.
The Moser-Scherck-Kemperman-Wehn Theorem is a basic tool in the proof by Gao
that a sequence of elements of an abelian  group $G$ with length
$|G|+d(G)$ contains a $|G|$--sub-sequence summing to $0,$ where $d(G)$ is the maximal size of a sequence of elements of $G$ having no non-empty zero-sum subsequence \cite{gaotnd}. A recent generalization of Gao Theorem is contained in \cite{HweightD}. Notice that the Moser-Scherck-Kemperman-Wehn Theorem
implies easily the following result of Shepherdson \cite{sheph}:

 For every  nonempty subset $S$ of a finite abelian group $G,$
there are $s_1, \cdots , s_k\in S$ such that $k\leq \lceil \frac{|G|}{|S|}\rceil $ and  $\sum _{1\leq i \leq k} s_i=0.$

Some of the above results are related to questions investigated in Combinatorics.
Let $\Gamma=(V,E)$  be a loopless finite (directed) graph  with  $\min \{|\Gamma (x)| : x\in V\}\ge 1$. It is well known that $D$ contains a directed cycle. The smallest cardinality  of such a cycle is called the girth of $D$ and will be denoted by $g(D)$.
In 1978,  Caccetta and H\"aggkvist \cite{CH}  conjectured that
 $$|V|\ge \min (d^+x: x\in V)(g(D)-1)+1.$$

 This conjecture is still largely open. The reader may find references and  results about this question in \cite{Bondy}.

 This conjecture were proved by the author for graphs with a transitive group of automorphisms in  \cite{HEJC}. This result applied to Cayley graphs shows the validity of Shepherdson's  result for all finite groups. Unfortunately we were not aware at that moment of Shepherdson's result. Our proof  \cite{HEJC} is based on the properties of atoms of a finite graph \cite{HATOM} and the Dirac-Menger's Theorem \cite{Bondy2,Nat,tv}.
 Another proof of the
 Caccetta and H\"aggkvist Conjecture for the last class of graphs,  based
 on  the Moser-Scherck-Kemperman-Wehn Theorem  and the representation of point-transitive graphs as coset graphs, was given by Nathanson  in \cite{Nat0}.
 A third proof, using atoms but avoiding the Dirac-Menger's  was given later by the author in \cite{Hspheres}.

In an e-mail cited by Mader \cite{mader},   Seymour suggested  the possibility of the existence of  $\min \{|\Gamma (x)| : x\in V\}\ge 1$ directed cycles  having pairwise a fixed vertex in the intersection.
 Such a property implies clearly the Caccetta and H\"aggkvist \cite{CH}  Conjecture.
 Mader obtained a positive answer
if $\Gamma $ has a transitive group of automorphisms\cite{mader}.

Generalizing all the above results, we prove the following:

Let $\Gamma =(V,E)$  be a  reflexive
relation having a transitive group of automorphisms. Let $v\in V$
and let $F$  be a  subset of $V$ with $F\cap \Gamma ^-(v)=\{v\}$. Then
\begin{itemize}
  \item[(i)] If $F$  is finite, then
$| \Gamma (F)\setminus F|\ge
|\Gamma (v)|-1.$
\item [(ii)]If $F$  is cofinite, then
$| \Gamma (F)\setminus F|\ge
|\Gamma ^- (v)|-1.$
\end{itemize}

In particular, for  i
  an integer $j\ge 1$ with  $\Gamma ^{j-1}(v)\cap \Gamma ^{-}(v)=\{v\}$, we have
 $
|\Gamma ^{j}  (v)|\ge | \Gamma ^{j-1}  (v)| +
|\Gamma (v)|-1.$

The finite case of the  last result is proved by Mader in \cite{mader}.
The same conclusion was obtain under the stronger hypothesis $\Gamma ^j(v)\cap \Gamma ^{-}(v)=\{v\}$ in \cite{Hspheres}. This weak form is enough to show Caccetta and H\"aggkvist Conjecture.

Let $G$  be group and let $B$ be a finite subset of $G.$  Applying this result to the Cayley graph  defined on $G$ by $B,$
we obtain Moser-Scherck-Kemperman-Wehn Theorem and its extension
to cofinite subsets:

Let $F$ be a finite or a cofinite subset of $G$ such that  $F\cap  B^{-1}=\{1\}$.
Then
$| (FB)\setminus F|\ge
|B|-1.$

Our extension to the cofinite case allows a very short proof for the
following result, proved by Losonczy \cite{los} in the abelian case and by    Eliahou-Lecouvey \cite{el} in the general case:

If $1\notin B,$ then here is a permutation $\sigma$ of $B$ such that $x\sigma (x)\notin B$ for every
$x\in B.$

We conclude the paper by a simple proof of Mader's result. According to our approach, Seymour's question may be formulated as follows:

Under which condition, a graph
satisfies the Moser-Scherck-Kemperman-Wehn property?

Our proofs  require properties, developed in Sections 3 and 4, of  weak atoms
and Moser sets. The proofs we obtain for these properties  are easier  than the proofs of the corresponding properties of the  atoms \cite{HATOM}.

The present work is essentially self-contained. Our proof of Mader's Theorem
requires the Dirac-Menger's Theorem. Also our proof of a result due  Losonczy and    Eliahou-Lecouvey requires K\"onig-Hall's Theorem.  Let us state below these two results:

\begin{theirtheorem} ( The Dirac-Menger's Theorem  \cite{Bondy2}) \label{menger}

Let $\Gamma=(V,E)$ be a finite reflexive graph Let $k$ be a
nonnegative integer. Let $x,y\in V$ such that $y\notin \Gamma (x).$ If $|\G (X)|\ge k$ for every subset $X$ with $x\in X$ and   $y\notin \Gamma (X)$,
then there are $k$  disjoint directed paths from $\Gamma (x)$ to $\Gamma ^-(y)$.
\end{theirtheorem}

\begin{theirtheorem} \label{kh} (K\"onig-Hall's Theorem \cite{Bondy2})
Let ${\Phi}\subset V\times W$ be a relation
with $|V|=|W|<\infty$. Then  the following conditions  are equivalent:
 \begin{itemize}
   \item There is a bijection $\phi :V\mapsto W$ with $\phi (x)\in {\Phi} (x),$ for every $x\in V.$
   \item  $|{\Phi}(Y)|\ge |Y|,$ for every  subset $Y$ of $V.$
 \end{itemize}

\end{theirtheorem}

The reader may find proofs of the Dirac-Menger's Theorem in the text books \cite{Nat,tv} and some applications of this result to Additive Number Theory. Notice that the K\"onig-Hall's Theorem follows by applying
the Dirac-Menger's Theorem to the graph obtained from $\Phi$ by adding two distinct vertices one dominating $V$ and the other dominated par $W.$
\section{Some Terminology}

Let $\Gamma =(V,E)$ be a  relation and let $X\subset V.$
The subrelation $\Gamma [X]$ induced on $X$ is by definition
$(X,(X\times X)\cap E).$
 A function $f : V  \longrightarrow V$ will be called a {\em homomorphism }if for all $x\in V$, we have
$\Gamma (f(x))=f(\Gamma (x))$. A bijective homomorphism is called an {\em automorphism}. The relation $\Gamma$ will be called {\em locally-finite}  if for all $x\in V,$ $|\Gamma (x)|$ and $|\Gamma ^-(x)|$ are finite.

The relation $\Gamma$ will be called {\em point-transitive}  if for all $x,y\in V,$ there is an automorphism $f$
such that $y=f(x)$.
Clearly a point-transitive relation is regular. Let  $S$  be a
subset of   $G$. The relation $(G,E),$ where  $ E=\{ (x,y) : x^{-1}y \
\in S \}$ is called a {\it Cayley relation}.  It will  be denoted by
$\mbox{Cay} (G,S)$.

Let $\Gamma =\mbox{Cay} (G,S)$   and  let   $F \subset G $.
Clearly
 $\Gamma (F)=FS ,$ where $FS=\{xy: x\in F \ \mbox{and}\ y\in S\}$.
 Cayley  graphs are clearly point-transitive.

 We identify graphs (directed graphs) and their relations.
The reader may replace everywhere  the term "relation" by "graph".

 We shall write
 $$\partial _{\Gamma}(X)= \Gamma (X)\setminus X .$$ We also write
 $$\nabla _{\Gamma}(X)= V\setminus \Gamma (X) .$$

When the context is clear the reference to $\Gamma$ will be omitted.

Let $F$  be a subset of  $V.$
Clearly $V=F\cup \nabla(F)\cup \partial (F)$ is a partition.
Notice that  $\partial^{-}(\nabla(F))\cap F=\emptyset,$ otherwise there exist $z\in F$ such that  $\Gamma (z)\cap (\nabla(F))\neq \emptyset$, a contradiction. Hence
\begin{equation}\label{id0}
\partial^{-}(\nabla(F))\subset \partial (F).\end{equation}

The last observation, used extensively in the isoperimetric method \cite{Hast}, contains a useful duality. The next lemma contains a useful sub-modular relation:

\begin{lemma} \cite{Hast}\label{partialsub}{Let $\Gamma =(V,E)$ be a  locally finite   reflexive
graph. Let   $X,Y$     be finite  subsets. Then

 \begin{equation}\label{submodularity}
|\partial  (X \cup Y)|+|\partial  (X \cap Y)|\le |\partial  (X )|+|\partial  ( Y)|
\end{equation}}
\end{lemma}

\begin{proof}

 Observe that
\begin{eqnarray*}
|\Gamma (X\cup Y)|&=&|\Gamma (X)\cup \Gamma (Y)|\\
&=&|\Gamma (X)|+|\Gamma (Y)|-|\Gamma(X)\cap \Gamma (Y)|\\
&\le& |\Gamma (X)|+|\Gamma (Y)|-|\Gamma(X\cap Y)|
\end{eqnarray*}
The result follows now by subtracting the equation $|X\cup Y|=|X|+|Y|-|X\cap Y|$.
\end{proof}

\section{Weak atoms}
 The reader  interested only in the finite case may ignore this section.
 Let $\Gamma =(V,E)$ be a locally finite  reflexive relation.
 We define the {\em weak connectivity} of $\Gamma$ as
\begin{equation}  \label{eq:kappa}
\kappa  (\Gamma)=\min  \{|\partial (X)|\  : \ \ 1\le |X|<\infty \}.
\end{equation}

 A subset $X$ achieving the  minimum in  (\ref{eq:kappa}) is called a
{\em weak fragment} of $\Gamma$. A weak fragment with minimum cardinality
 will be called a {\em weak atom}.

\begin{proposition}  {
Let $\Gamma =(V,E)$ be a
 locally finite   relation. If $F_1$ and $F_2$ are weak
 fragments with a nonempty intersection, then $F_1\cup F_2$ and $F_1\cap F_2$ are weak fragments. In particular, two distinct weak atoms are disjoint. Assume furthermore that $\Gamma$ is point--transitive and
let $A$ be an atom of $\Gamma$. Then the weak atoms induce a partition of $V$. Moreover the subrelation $\Gamma [A]$ induced on $A$ is a point-transitive relation, for any weak atom $A$.
\label{atom}} \end{proposition}

\begin{proof}
By (\ref{submodularity}),
$2\kappa (\Gamma)\le |\partial  (F_1 \cap F_2)|+|\partial  (F_1 \cup F_2)|\le |\partial  (F_1 )|+|\partial ( F_2)|\le 2\kappa (\Gamma).$
Let $A$ be a weak atom and take $v\in A$. For every $y\in V,$ there is an automorphism $\phi$ such that $\phi (v)=y.$ Hence $y$ belongs to the atom $\phi (A).$ In particular, the set of atoms is a partition of $V.$ If $y\in A,$ then $\phi(A)=A,$ and therefore $\phi/A$ is an automorphism of $\Gamma[A]$ with $\phi/A(v)=y$. Thus $\Gamma [A]$  is a point-transitive relation.
\end{proof}

\section{Moser sets}
The formalism we develop here is motivated from one side by the Moser-Scherck-Kempermann-Wehn Theorem and by the notion of vertex-fragment
(a possibly empty set) in finite graphs considered by Mader in \cite{mader}  from the other side. Our Moser sets are related to Mader's
vertex-fragments. But our concept is never empty, works in the infinite case too and leads to easier proofs. In particular, we do not need a duality between positive and negative vertex-fragments, in the spirit of the one introduced in \cite{HATOM}, used extensively in Mader arguments. Also the Dirac-Menger's Theorem,
present in beginning of Mader's formalism, is not needed to prove our main result, generalizing of the Moser-Scherck-Kempermann-Wehn Theorem. We use it only to give a simple proof of Mader's result.

 Let $\Gamma =(V,E)$ be a locally finite  reflexive relation and let $v\in V$.
A set $F$ is said to be a $v$--Moser set if $\Gamma ^-(v)\cap F=\{v\}.$ Put
$$\kappa _{\Gamma}(v)=\min \{|\partial (X)| : X \ \mbox{is a} \ v \mbox{--Moser set}\}.$$
A Moser set $X$ with $|\partial (X)|=\kappa _v(\Gamma)$ will be called a $v$--fragment.

\begin{lemma} Let $F_1$ and $F_2$ be two $v$--fragments.
Then $F_1\cap F_2$ and  $F_1\cup F_2$ are $v$--fragments.
In particular, there exists a $v$--fragment $K_v$ contained in  every $v$--fragment.
\end{lemma}

\begin{proof}
Notice that the intersection
and union of two $v$--Moser sets are $v$--Moser sets.
By (\ref{submodularity}),
$2\kappa _{\Gamma}(v) \le |\partial  (F_1 \cap F_2)|+|\partial  (F_1 \cup F_2)|\le |\partial  (F_1 )|+|\partial ( F_2)|\le 2\kappa _{\Gamma}(v).$
\end{proof}

We shall denote the minimal $v$--fragment containing  $v$ by $K_v$.

Consider the subgraph $\Theta =(V,F),$ where $F=\{(x,y)\in E :  y\in K_x\}$.
 Set $\Psi (x) =\{y :  K_y \subset K_x \ \mbox{and}\  y\in \Theta (x)\}$.

\begin{lemma} \label{mader}
$\Theta^-(x)\setminus \Psi (x) \subset \G(K_x)\setminus \Gamma (x)$.
\end{lemma}

\begin{proof}

Take $y\in \Theta^-(x)\cap (V\setminus \Gamma (K_x))$. Since   $x\in K_y,$ we have by (\ref{id0}), $\Gamma ^-(y)\subset
\Gamma ^-(V\setminus \Gamma (K_x))\subset V\setminus K_x$. It follows that  $\Gamma ^-(y)\cap (K_x \cup K_y)=\Gamma ^-(y)\cap  K_y=\{y\}.$ Thus $K_x \cup K_y$ is a $y$--Moser set.
Noticing that $K_x \cap K_y$ is an $x$--Moser set,  we have  by (\ref{submodularity}),

$\kappa _{\Gamma}(y)+ \kappa _{\Gamma} (x)\le |\partial  (K_x \cap K_y)|+|\partial  (K_x \cup K_y)|\le |\partial  (K_x )|+|\partial ( K_y)|\le \kappa _{\Gamma}(x)+ \kappa _{\Gamma} (y).$

It follows that $K_x \cap K_y$ is a $x$--Moser set and hence $K_x=K_x \cap K_y$. Thus $y\in \Psi(x).$ Therefore $$\Theta^-(x)\cap (V\setminus \Gamma (K_x))\subset \Psi (x).$$

Take  $y\in \Theta^-(x)\setminus \Psi (x)$. Then $x\in \Theta (y)\subset \Gamma (y).$  By the last relation, we must have $y\in \Gamma (K_x)$.
Since $K_x\cap \Gamma ^-(x)=\{x\},$ we have $y\notin
\Gamma (x)$. In particular $\Theta^-(x)\setminus \Psi (x)\subset \Gamma(x)$.
   The lemma follows now.
\end{proof}

\section{The result and its applications}

\begin{theorem}\label{main}
Let $\Gamma =(V,E)$  be a point-transitive reflexive locally finite
relation and  let $v\in V.$
Let $F$  be a  subset of $V$ with $F\cap \Gamma ^-(v)=\{v\}$.
\begin{itemize}
  \item[(i)] If $F$  is finite, then
$| \Gamma (F)\setminus F|\ge
|\Gamma (v)|-1.$
\item [(ii)]If $F$  is cofinite, then
$| \Gamma (F)\setminus F|\ge
|\Gamma ^- (v)|-1.$
\end{itemize}

\end{theorem}
\begin{proof}
Recall that for every $x\in V,$ $|\Gamma (v)|=|\Gamma (x)|$. This is an immediate consequence
of the transitive action of automorphism group. Moreover if $V$ is finite,
$|V||\Gamma (v)|=|E|=|V||\Gamma^- (v)|$.
Thus $|\Gamma (v)|=|\Gamma ^- (v)|.$  In particular, (i) and (ii) are equivalent in the finite case.
For every $x\in V,$ there is an automorphism $\phi$ such that $\phi (v)=x.$ Clearly
$\phi (K_v)= \phi (K_x)$ and hence
$|K_v|=|K_x|.$ If $x\in \Gamma (v),$  then $v\in K_v\setminus K_x,$ and thus we have $K_v \neq K_x$ and hence $K_v\not\subset
K_x$. It follows that $\Psi (x)=\{x\}.$

{\bf Case 1} $V$ is finite.
Since $\sum _{x\in V}|\Psi ^{-}(x)|=\sum _{x\in V}|\Psi (x)|,$
there is a $u\in V$ with $|\Theta(u)|\le  |\Theta^-(u)|$.

 By Lemma \ref{mader},
 $\Theta^-(u) \setminus \{u\} =\Theta^-(u)\setminus \Psi (u)\subset \G(K_u)\setminus \Gamma (u)$. Now we have
\begin{eqnarray*}|\G(K_u)| &\ge& |\Theta ^-(u)\setminus \{u\} |+| \Gamma (u) \setminus K_u|\\&\ge& |\Theta(u)\setminus \{u\}|
+| \Gamma (u)\setminus K_u|\\&\ge& |(\Gamma(u)\cap K_u)\setminus \{u\}|
+| \Gamma (u)\setminus K_u|= |\Gamma (u)|-1.\end{eqnarray*}
Since $ \{u\}$ is a Moser set, we have $|K_u|=1$. It follows that $|K_v|=|K_u|=1.$
Since $F$ is a Moser set, we have $$
| \Gamma (F)\setminus F|\ge  \kappa _v (\Gamma )=
|\Gamma (K_v)|-1=|\Gamma (v)|-1.$$
 The result follows in this case.

{\bf Case 2} $V$ is infinite.

{\bf Subcase 2.1} $F$ is finite.

Let $A$ be an atom of $\Gamma$ containing $v.$ Assume first that $\kappa (\Gamma)=0.$
Every point  $x\in V\setminus A$ is contained in an atom $A_x,$ disjoint from $A$ by Proposition \ref{atom} with
$\partial (A_x)|=0.$ It follows that $\Gamma (F\setminus v)\cap A=\emptyset$. Since
$A$ is finite the result holds for $A$ by Case 1. It follows that $|\Gamma (F)|=|\Gamma (F\setminus A)|+
|\Gamma (F\cap A)|\ge |F\setminus A|+|F\cap A|+|\Gamma (v)|-1.$ So we may assume that $\kappa (\Gamma)>0.$
Observe that $F\cap A$ is a $v$--Moser set of $A$. By case 1, $$
| \Gamma (F\cap A)|\ge |A\cap F|+|A\cap \Gamma (v)|-1.$$
 By the definition of $\kappa,$ we have $|\partial (F\cup,A)|\ge \kappa=\partial (A).$
Notice that $\partial (F\cup A)\setminus \partial (F)\subset \partial (A) \setminus \Gamma (v)$,
and hence \begin{eqnarray*}|\partial (F)|&\ge& | \G (F)\cap A|+ |\partial (F\cap A)\cap \partial (F)|\\ &\ge&
| \G (F)\cap A|+ |\partial (F\cup A)|-|\partial (A) \setminus
 \Gamma (v)|\\&\ge& |A\cap \Gamma (v)|-1+|\partial (A)|-|\partial (A)\setminus \Gamma (v)|\\&=& |A\cap \Gamma (v)|-1+|\partial (A)\cap \Gamma (v)|=|\Gamma (v)|-1.\end{eqnarray*}

 {\bf Subcase 2.2} $F$ is cofinite.
 Put $R=(V\setminus \Gamma (F))\cup \{v\}.$ It follows that $R\cap \Gamma (v)=\{v\}\cap \Gamma (v)=\{v\}.$ In particular, $R$ is a $v$--Moser set for the relation
 $\Gamma ^-$. Also we have using (\ref{id0}), $\G ^-(R)\subset \G^-(V\setminus \Gamma (F))\cup (\G^-(v)\setminus R)\subset \G(F),$ since $F\cap \Gamma ^-(v)=\{v\}.$

 By Subcase 2.1, $|\G (F)|\ge |\G^-(V\setminus \Gamma (F))|\ge
 |\G^-(R)|\ge |\Gamma^- (v)|-1.$
\end{proof}

\begin{corollary}\label{sphere}
Let $\Gamma =(V,E)$  be a point-symmetric reflexive locally finite relation and  let $v\in V.$
Let $j\ge 1$ be an integer such that  $\Gamma ^{j-1}(v)\cap \Gamma ^{-}(v)=\{v\}$. Then
 $$
|\Gamma ^{j}  (v)|\ge | \Gamma ^{j-1}  (v)| +
|\Gamma (v)|-1.$$
\end{corollary}

The proof is an immediate consequence of Theorem \ref{main}.

The finite case of the  corollary is proved by Mader in \cite{mader}.
The same conclusion was obtained under the stronger hypothesis $\Gamma ^j(v)\cap \Gamma ^{-}(v)=\{v\}$ in \cite{Hspheres}. This weak form is enough to show Caccetta and H\"aggkvist Conjecture.

\begin{corollary}\label{mskw} (The Moser-Scherck-Kemperman-Wehn Theorem and its extension to cofinite sets)

Let $G$  be group and let $S$ be a finite subset of $G.$ Let $F$ be a finite or a cofinite subset of $G$ such that  $F\cap  S^{-1}=\{1\}$.
Then
$| (FS)\setminus F|\ge
|S|-1.$

\end{corollary}
\begin{proof}
The corollary follows by applying Theorem \ref{main} to the Cayley relation
defined on $G$ by $S$.\end{proof}

%subsection{Terminology}
Let $G$ be a group and let  $X\subset G.$ We shall write $\overline{X}=G\setminus X$ and
$\tilde{X}=X\cup \{1\}.$

\begin{corollary}\label{ce} (Losonczy in the abelian case,   Eliahou-Lecouvey in the general case)

Let $G$  be group and let $A$ be a finite subset of $G\setminus \{1\}.$  There is a permutation $\sigma$ of $S$ such that $x\sigma (x)\notin A$ for every
$x\in A.$

\end{corollary}
\begin{proof}
Define a relation $\Phi$ on $ A,$ where $\Phi (x)=(x^{-1}\oa)\cap A.$
For every $B\subset A^{-1},$ we have clearly $\oa\cap \tilde{B}=\{1\}.$
Since $\oa$ is a cofinite subset, we have by Corollary \ref{mskw},
$|B|\le |(\oa \tilde{B})\setminus \oa|=|(\oa B)\setminus \oa|=|\Phi (B^{-1})|.$
The existence $\sigma$ follows now by Theorem \ref{kh}.
\end{proof}

\begin{corollary}(Mader \cite{mader})\label{maderth}
Let $\Gamma =(V,E)$  be a point-transitive loopless  finite
relation and  let $v\in V.$ There exist $|\Gamma (v)| $ directed cycles in a with a  pairwise intersection $=\{v\}$.
\end{corollary}
\begin{proof} Let $\Phi $ be the  graph obtained by a adding vertex $  v'\notin V,$ with $\Phi(v')=\Gamma ^- (v).$ Let $\Theta $ be the reflexive
closure $\Phi$ (obtained by adding loops everywhere. Let $F$ be a subset of $V$
with $v\in F$ and $v'\notin \Theta (F).$ Then clearly $F$ is a $v$--Moser set.
By Theorem \ref{main}, $|\G(F)|\ge |\Theta (v)|.$ By the Dirac-Menger's Theorem \ref{menger},
$\Phi$ contains $|\Gamma (v)|$ disjoint paths from $\Gamma (v)$ to $\Gamma ^-(v')=\Gamma ^-(v)$.
Adding $v$ to each of these paths, we see the existence of $|\Gamma (v)|$ cycles verifying the Corollary.
\end{proof}

%\end{document}

\end{document}